OpenStudy (anonymous):

(3a^2)^3 idk if I did this right is the answer 27a^6 or 27a^5

OpenStudy (solomonzelman):

you got the 27 right, and a is raised to what power ?

OpenStudy (anonymous):

i forget if it is add the exponents or times them

OpenStudy (solomonzelman):

RULE: \(\large\color{blue}{ (a^b)^c=a^{b\times c}}\) (But, \(\large\color{black}{ (a^b)(a^c)=a^{b+ c}}\) )

OpenStudy (solomonzelman):

so ti would be to the exponent of \(\large\color{black}{ ^{2\times 3}}\)

OpenStudy (anonymous):

okay so its 27^6 then right?

OpenStudy (solomonzelman):

yes

OpenStudy (solomonzelman):

sorry for the late reply (due to a chess game)